# FlightPatchNet: Multi-Scale Patch Network with Differential Coding for Flight Trajectory Prediction

### v1

@@ 4.5 Ablation Study
Table 4: Performance comparisons on ablative variants. The best results are highlighted in bold.
Case Horizon Lon() Lat() Alt(m)
1 3 9 15 1 3 9 15 1 3 9 15
w/o global temporal attention MAE 0.00051 0.00190 0.00667 0.01232 0.00034 0.00132 0.00466 0.00876 20.59 28.15 66.62 112.43
RMSE 0.00090 0.00308 0.01085 0.02005 0.00066 0.00222 0.00791 0.01486 136.79 112.10 163.78 220.83
w/o scale fusion MAE 0.00053 0.00169 0.00609 0.01112 0.00035 0.00114 0.00409 0.00759 24.17 32.70 91.06 162.20
RMSE 0.00092 0.00268 0.00975 0.01787 0.00067 0.00188 0.00688 0.01280 130.05 99.54 194.25 282.07
w/o channel fusion MAE 0.00050 0.00166 0.00573 0.01059 0.00034 0.00112 0.00398 0.00727 20.04 29.22 73.23 131.98
RMSE 0.00089 0.00265 0.00924 0.01707 0.00065 0.00187 0.00667 0.01240 159.47 122.32 174.36 250.07
FlightPatchNet MAE 0.00048 0.00153 0.00546 0.00966 0.00032 0.00105 0.00381 0.00678 13.34 32.65 78.57 123.97
RMSE 0.00087 0.00233 0.00885 0.01577 0.00064 0.00175 0.00652 0.01174 123.78 121.48 174.63 244.34
@@
5 Conclusion
In this paper, we propose FlightPatchNet, a multi-scale patch network with differential coding for short-term FTP task. The differential coding is leveraged to reduce the significant differences in original data range and reflect the temporal variations in realistic flight trajectories. The multi-scale patch network is designed to explore global trends and local details based on divided patches of different sizes, and integrate scale-wise correlations and inter-variable relationships for complete temporal modeling. Extensive experiments on a real-world dataset demonstrate that FlightPatchNet achieves the most competitive performance.
@@ A.1 Dataset Preprocessing and Description
(3) Format Transformation: We transform the horizontal velocity into $V_{x}$ and $V_{y}$ according to the angle, where $V_{x}$ is the velocity in the longitude dimension and $V_{y}$ is the velocity in the latitude dimension. In this way, the features become longitude, latitude, altitude, $V_{x}$, $V_{y}$ and $V_{z}$.
@@
After the above preprocessing, 274,605 flight trajectories are selected into our dataset. The range of longitude, latitude and altitude are $[-179.86396^{\circ},178.82147^{\circ}]$, $[-46.42435^{\circ},70.32590^{\circ}]$ and $[0,21031.00m]$, respectively. The interval between two adjacent flight trajectory points is 10 seconds.
@@ A.2 Baseline Methods
We briefly describe the selected 5 competitive baselines as follows:
  • LSTM [19]: Based on two layers of LSTM (with 30 and 60 nodes respectively) to encode each trajectory point, and future trajectories are predicted through a fully connected layer.
  • Bi-LSTM [32]: Based on two layers of Bi-LSTM (with 200 and 50 nodes respectively) to encode each trajectory point, and future trajectories are predicted through a fully connected layer.
  • CNN-LSTM [18]: Based on two layers of one-dimensional CNN (the convolution kernel size is $1\times 3$) and two layers of LSTM (with 50 nodes) to encode each trajectory point, and future trajectories are predicted through a fully connected layer.
  • FlightBERT [15]: It utilizes a BE representation to convert the scalar attributes of the flight trajectory into binary vectors, considering the FTP task as a multi binary classification problem. It uses 18, 16, 11 and 11 bits to encode the real values (decimals) of longitude, latitude, altitude and velocities into BE representation respectively.
  • FlightBERT++ [16]: It inherits the BE representation from the FlightBERT and introduces a differential prediction paradigm, which aims to predict the differential values of the trajectory attributes instead of the absolute values.
@@ A.3 Implementation Details
For fairness, all the models follow the same experimental setup with look-back window $L=60$ and prediction horizon $T\in\{1,3,9,15\}$, which means the observation time is 10 minutes and the forecasting time is 10 seconds, 30 seconds, 1.5 minutes, 2.5 minutes. The patch sizes in multi-scale patch mixer blocks are set to {30, 20, 10, 6, 2}. The dimension of temporal embedding $d$ is 128. For all the MSA in this paper, the head number is 8 and the attention layer $l$ is 3. The learning rate is set as $10^{-4}$ for all experiments. Our method is trained with MSE loss, using the Adam optimizer [40]. The training process is early stopped within 30 epochs. The training would be terminated early if the validation loss does not decrease for three consecutive rounds. The model is implemented in PyTorch 2.2.1 and trained on a single NVIDIA RTX 4080 GPU with 16GB memory.
@@ B.1 Hyper-Parameter Sensitivity
We perform experiments on the different number of scales and report the MAE and RMSE results. As shown in Figure 4, we can observe that when the number of scales increases from 2 to 5, the performance of FlightPatchNet is constantly improved. This is because FlightPatchNet can capture diverse global and local temporal patterns under different scales. When the number of scales increases up to 6, the performance starts to deteriorate. This indicates that a certain number of scales is sufficient for temporal modeling, and excessive scales may lead to the overfitting problem.
Figure 4: MAE and RMSE with different number of scales for prediction horizon $T\in\{1,3,9,15\}$
@@
Figure 6: MAE and RMSE of different look-back window sizes for prediction horizon $T=15$
@@
patch sizes Lon() Lat() Alt(m)
@@
2,6,10,20,30 $\uparrow$ MAE 0.00098 0.00155 0.00548 0.01008 0.00099 0.00106 0.00385 0.00697 54.02 33.47 79.39 127.51
RMSE 0.00187 0.00241 0.00887 0.01642 0.00131 0.00183 0.00656 0.01197 81.92 110.68 184.54 248.16
$\downarrow$ MAE 0.00048 0.00153 0.00546 0.00966 0.00032 0.00105 0.00381 0.00678 13.34 32.65 78.57 123.97
RMSE 0.00087 0.00233 0.00885 0.01577 0.00064 0.00175 0.00652 0.01174 129.65 121.78 174.63 244.34
3,4,6,20,40 $\uparrow$ MAE 0.00098 0.00155 0.00556 0.00997 0.00064 0.00105 0.00383 0.00704 39.77 32.15 76.38 124.18
RMSE 0.00188 0.00247 0.00901 0.01631 0.00131 0.00175 0.00655 0.01210 64.86 124.20 177.46 243.46
$\downarrow$ MAE 0.00097 0.00153 0.00542 0.00963 0.00063 0.00104 0.00369 0.00670 43.46 28.96 79.27 128.13
RMSE 0.00187 0.00245 0.00879 0.01582 0.00130 0.00174 0.00631 0.01167 64.50 115.76 176.96 251.72
3,6,40 $\uparrow$ MAE 0.00048 0.00156 0.00536 0.00994 0.00035 0.00105 0.00370 0.00691 14.96 31.42 79.07 117.43
RMSE 0.00087 0.00248 0.00876 0.01628 0.00065 0.00176 0.00634 0.01193 107.43 118.36 177.40 238.87
$\downarrow$ MAE 0.00048 0.00153 0.00534 0.00988 0.00033 0.00103 0.00368 0.00685 16.81 31.26 71.96 118.66
RMSE 0.00087 0.00244 0.00870 0.01620 0.00064 0.00173 0.00633 0.01186 145.25 114.33 175.63 236.32
@@ B.2 Error Bar
Model Horizon Lon() Lat() Alt(m)
@@
FlightBERT++ 1 0.00173$\pm$6.45e-5 0.00360$\pm$8.28e-5 0.00085$\pm$3.33e-5 0.00148$\pm$1.32e-4 9.39$\pm$1.79 175.29$\pm$29.09
3 0.00317$\pm$2.61e-4 0.00659$\pm$4.35e-5 0.00210$\pm$3.05e-4 0.00425$\pm$1.24e-4 21.89$\pm$5.58 167.16$\pm$46.39
9 0.00871$\pm$1.74e-4 0.01846$\pm$4.45e-4 0.00612$\pm$6.07e-4 0.00959$\pm$2.19e-4 47.84$\pm$2.87 327.93$\pm$52.84
15 0.01187$\pm$5.91e-5 0.03131$\pm$5.33e-4 0.01048$\pm$3.69e-4 0.02127$\pm$2.01e-4 78.46$\pm$8.13 384.18$\pm$51.82
FlightPatchNet (Ours) 1 0.00048$\pm$1.24e-5 0.00087$\pm$1.02e-5 0.00032$\pm$1.06e-5 0.00064$\pm$8.45e-6 13.34$\pm$9.43 123.78$\pm$15.13
3 0.00153$\pm$3.19e-5 0.00233$\pm$5.44e-4 0.00105$\pm$1.19e-5 0.00175$\pm$2.36e-5 32.65$\pm$1.76 121.48$\pm$2.81
9 0.00546$\pm$1.54e-4 0.00885$\pm$2.16e-4 0.00381$\pm$7.47e-5 0.00652$\pm$7.76e-5 78.57$\pm$2.66 174.63$\pm$6.87
15 0.00966$\pm$3.65e-4 0.01577$\pm$5.49e-4 0.00678$\pm$2.58e-4 0.01174$\pm$3.53e-4 123.97$\pm$5.72 244.34$\pm$6.91
B.3 Model Complexity
As shown in Table 7, our proposed FlightPatchNet achieves the greatest efficiency and has relatively small parameters compared to other models. For multi-step prediction, the DMS-based models (FlightPatchNet, FlightBERT++) demonstrate significant improvements in computational performance compared to the IMS-based models (FlightBERT, LSTM, Bi-LSTM, CNN-LSTM). In addition, FlightPatchNet is lightweight compared to FlightBERT++ and FlightBERT, which indicates our model can provide a reliable solution for real-time air transportation management.
Models Parameters (MB) FLOPs (M) Running Time (s/iter)
FlightPatchNet(ours) 5.69 64.38 0.0069
FlightBERT++ 44.26 3000 0.0112
FlightBERT 25.31 1620 0.2406
LSTM 0.03 1.67 0.0583
Bi-LSTM 0.51 31.15 0.1241
CNN-LSTM 0.04 1.22 0.0429
Appendix C Visualization
Visualization of FlightPatchNet Predictions
Figure 7: Visualization of the ground truth and predictions of FlightPatchNet. The series of altitude are in original data while those of longitude and latitude are in differential data, all denoted by meters.
Visualization of FlightPatchNet Predictions for Altitude
We present the visualization of FlightPatchNet predictions and ground truth for altitude in Figure 8. As shown in Figure 8, when the series of altitude are relatively smooth and stationary with obvious global trends, FlightPatchNet can effectively capture these trends and make accurate predictions. When the series suffers from many change points caused by frequent abrupt fluctuations, as depicted in Figure 8 and Figure 8, FlightPatchNet tends to focus on the irregular change points during prediction, leading to a large deviation from the ground truth. As a result, FlightPatchNet struggles to capture the real temporal variations in altitude and fails to provide accurate predictions.
Figure 8: Visualization of predictions of FlightPatchNet and ground truth for altitude.
3D Trajectory Visualization
@@ Appendix C Visualization
Figure 9: Visualization of flight trajectory prediction results when the prediction horizon $T=15$. The look-back window size $L=60$.
Appendix D Limitation and Future Work
FlightPatchNet has shown the most competitive performance for flight trajectory prediction. It is easy to implement and has fewer parameters, presenting a promising solution for real-time air traffic control applications. However, we also acknowledge the limitations of our work. Since the original series of altitude contains many fluctuations caused by unexpected noise, our primary focus on modeling temporal variations has led to a large bias towards output predictions of altitude. In the future, we will further explore temporal modeling of altitude and consider graph networks to overcome the prediction bias affected by unexpected noise.

### v3

@@ 4.3 Ablation Study
Table 3: Performance comparisons on ablative variants. The best results are highlighted in bold. Hor represents the prediction horizon $T\in\{1,3,9,15\}$.
Case Hor
Lon
($0.00001^{\circ}$)
Lat
($0.00001^{\circ}$)
Alt
MAE RMSE MAE RMSE MAE RMSE
w/o
global
temporal
attention
1 51 90 34 66 21 137
3 190 308 132 222 28 112
9 667 1085 466 791 67 164
15 1232 2005 876 1486 112 221
w/o
scale
fusion
1 53 92 35 67 24 130
3 169 268 114 188 33 100
9 609 975 409 688 91 194
15 1112 1787 759 1280 162 282
w/o
channel
fusion
1 50 89 34 65 20 160
3 166 265 112 187 29 122
9 573 924 398 667 73 174
15 1059 1707 727 1240 132 250
FlightPatchNet 1 48 87 32 64 13 124
3 153 233 105 175 33 122
9 546 885 381 652 79 175
15 966 1577 678 1174 124 244
@@
4.4 Model Complexity
Type Models
Parameters
FLOPs
Running Time
(s/iter)
DMS FlightPatchNet 5.69 64.38 0.0069
FlightBERT++ 44.26 3000.00 0.0112
Pathformer 2.64 258.55 0.02174
TimeMixer 0.45 3115.00 0.0037
TimesNet 37.50 196159.87 0.0534
MICN 1.28 1235.83 0.0012
IMS FlightBERT 25.31 1620.00 0.2406
LSTM 0.03 1.67 0.0583
Bi-LSTM 0.51 31.15 0.1241
CNN-LSTM 0.04 1.22 0.0429
WTFTP 0.23 60.00 0.0290
As shown in Table 4, our proposed FlightPatchNet achieves the greatest efficiency and has relatively fewer parameters compared to other models. For multi-step prediction, the DMS-based models demonstrate significant improvements in computational performance compared to the IMS-based models. In addition, FlightPatchNet is lightweight compared to FlightBERT++ and FlightBERT, which indicates our model can provide a promising solution for real-time air transportation management.
5 Conclusion, Limitation and Future Work
Conclusion
In this paper, we propose FlightPatchNet, a multi-scale patch network with differential coding for the short-term FTP task. The differential coding is leveraged to reduce the significant differences in the original data range and reflect the temporal variations in realistic flight trajectories. The multi-scale patch network is designed to explore global trends and local details based on divided patches of different sizes, and integrate scale-wise correlations and inter-variable relationships for complete temporal modeling. Extensive experiments on a real-world dataset demonstrate that FlightPatchNet achieves the most competitive performance and offers a significant reduction in computational complexity, presenting a promising solution for real-time air traffic control applications.
Limitation and Future Work
The original series of altitude contains many unexpected noises. Our primary focus on modeling temporal variations enables FlightPatchNet to concentrate more on these irregular change points during altitude prediction, leading to a large bias. In the future, we will further explore temporal modeling of altitude and investigate robust noise-handling techniques such as moving average to make the series smoother and less sensitive to noise. In addition, the data-missing problem occurs commonly in realistic flight trajectory series, making downstream analysis difficult. Thus, we attempt to incorporate data imputation methods into our model to enhance the model applicability and provide a general framework for flight trajectory prediction.
Appendix
@@ A.1 Dataset Preprocessing and Description
(3) Velocity Transformation: We transform the horizontal velocity into $V_{x}$ and $V_{y}$ according to the angle, where $V_{x}$ is the velocity in the longitude dimension and $V_{y}$ is the velocity in the latitude dimension. In this way, the features become longitude, latitude, altitude, $V_{x}$, $V_{y}$ and $V_{z}$.
@@
After the above preprocessing, 274,605 flight trajectories are selected into our dataset. The range of longitude, latitude and altitude are $179.86396^{\circ},178.82147^{\circ}]$, $46.42435^{\circ},70.32590^{\circ}]$ and $[0,21031.00m]$, respectively. The interval between two adjacent flight trajectory points is 10 seconds.
@@ A.2 Baseline Methods
We briefly describe the selected 10 competitive baselines as follows:
  • LSTM [Shi et al., 2018]: Based on two layers of LSTM (with 30 and 60 nodes respectively) to encode each trajectory point, and future trajectories are predicted through a fully connected layer.
  • CNN-LSTM [Ma and Tian, 2020]: Based on two layers of one-dimensional CNN (the convolution kernel size is $1\times 3$) and two layers of LSTM (with 50 nodes) to encode each trajectory point, and future trajectories are predicted through a fully connected layer.
  • Bi-LSTM [Sahadevan et al., 2022]: Based on two layers of Bi-LSTM (with 200 and 50 nodes respectively) to encode each trajectory point, and future trajectories are predicted through a fully connected layer.
  • FlightBERT [Guo et al., 2023]: It utilizes a BE representation to convert the scalar attributes of the flight trajectory into binary vectors, considering the FTP task as a multi binary classification problem. It uses 18, 16, 11 and 11 bits to encode the real values (decimals) of longitude, latitude, altitude and velocities into BE representation respectively.
  • FlightBERT++ [Guo et al., 2024]: It inherits the BE representation from the FlightBERT and introduces a differential prediction paradigm, which aims to predict the differential values of the trajectory attributes instead of the absolute values.
  • WTFTP [Zhang et al., 2023]: It is an IMS-based method that utilizes discrete wavelet transform (DWT) to decompose the input flight trajectory into wavelet coefficients and predicts future trajectories based on the generated wavelet coefficients by inverse discrete wavelet transform (IDWT).
  • TimeMixer [Wang et al., 2024]: It is a fully MLP-based architecture that mixes the decomposed seasonal and trend components in fine-to-coarse and coarse-to-fine directions separately and ensembles multiple predictors to utilize complementary forecasting capabilities in multi-scale observations.
  • TimesNet [Wu et al., 2023a]: It is a task-general foundational model for time series analysis, which disentangles complex temporal variations into multiple intra-period and inter-period variations. A parameter efficient inception block is employed to capture these temporal variations in 2D space.
  • MICN [Wang et al., 2023]: It adopts a multi-scale branch structure to capture the underlying information in time series. Downsampling one-dimensional convolution is used for local feature extraction and isometric convolution is employed for global correlation discovery.
  • Pathformer [Chen et al., 2024]: It is a multi-scale Transformer with adaptive pathways, which integrates both temporal resolution and temporal distance for multi-scale modeling. A multi-scale router with temporal decomposition and an aggregator work together to realize adaptive multi-scale modeling for time series.
@@ A.3 Implementation Details
For fairness, all the models follow the same experimental setup with look-back window $L=60$ and prediction horizon $T\in\{1,3,9,15\}$, which means the observation time is 10 minutes and the forecasting time is 10 seconds, 30 seconds, 1.5 minutes, 2.5 minutes. The patch sizes in multi-scale patch mixer blocks are set to {30, 20, 10, 6, 2}. The dimension of temporal embedding $d$ is 128. For all the MSA in this paper, the head number is 8 and the attention layer $l$ is 3. The learning rate is set as $10^{-4}$ for all experiments. Our method is trained with MSE loss, using the Adam optimizer. The training process is early stopped within 30 epochs. The training would be terminated early if the validation loss does not decrease for three consecutive rounds. The model is implemented in PyTorch 2.2.1 and trained on a single NVIDIA RTX 3090 GPU with 24GB memory.
@@ B.1 Hyper-Parameter Sensitivity
We perform experiments on different number of scales and report the MAE and RMSE results. As shown in Figure 9, we can observe that when the number of scales increases from 2 to 5, the performance of FlightPatchNet is constantly improved. This is because FlightPatchNet can capture diverse global and local temporal patterns under different scales. When the number of scales increases up to 6, the performance starts to deteriorate. This indicates that a certain number of scales is sufficient for temporal modeling, and excessive scales may lead to the overfitting problem.
Figure 9: MAE and RMSE with different number of scales for prediction horizon $T\in\{1,3,9,15\}$.
@@
Figure 11: MAE and RMSE of different look-back window sizes $L$ for prediction horizon $T=15$.
@@
patch sizes Lon($0.00001^{\circ}$) Lat($0.00001^{\circ}$) Alt(m)
@@
2,6,10,20,30 $\uparrow$ MAE 98 155 548 1008 99 106 385 697 54.02 33.47 79.39 127.51
RMSE 187 241 887 1642 131 183 656 1197 81.92 110.68 184.54 248.16
$\downarrow$ MAE 48 153 546 966 32 105 381 678 13.34 32.65 78.57 123.97
RMSE 87 233 885 1577 64 175 652 1174 129.65 121.78 174.63 244.34
3,4,6,20,40 $\uparrow$ MAE 98 155 556 997 64 105 383 704 39.77 32.15 76.38 124.18
RMSE 188 247 901 1631 131 175 655 1210 64.86 124.20 177.46 243.46
$\downarrow$ MAE 97 153 542 963 63 104 369 670 43.46 28.96 79.27 128.13
RMSE 187 245 879 1582 130 174 631 1167 64.50 115.76 176.96 251.72
3,6,40 $\uparrow$ MAE 48 156 536 994 35 105 370 691 14.96 31.42 79.07 117.43
RMSE 87 248 876 1628 65 176 634 1193 107.43 118.36 177.40 238.87
$\downarrow$ MAE 48 153 534 988 33 103 368 685 16.81 31.26 71.96 118.66
RMSE 87 244 870 1620 64 173 633 1186 145.25 114.33 175.63 236.32
@@ B.2 Error Bar
Model Horizon Lon($0.00001^{\circ}$) Lat($0.00001^{\circ}$) Alt(m)
@@
[…]
FlightPatchNet
[…]
Appendix C 3D Trajectory Visualization
@@
Figure 12: Visualization of flight trajectory prediction results when the prediction horizon $T=15$ and look-back window size $L=60$.